d^2-6dd'+9d'^2=12x^2+36xy
1
Expert's answer
2020-09-20T18:23:40-0400

Given differential equation is "(D^2-6DD'+9D'^2)z=12x^2+36xy"

where "D = \frac{d}{dx}, D' = \frac{d}{dy}" .

"\implies (D-3D')^2 z = 12 x^2 + 36 xy"

So, Auxiliary equation is "(D-3D')^2 = 0 \implies D-3D' = 0 , D - 3D' = 0" .

Hence, Complementary function (C.F.) = "\phi_1(y+3x) + x \phi_2(y+3x)" .

Particular Integral (P.I.) = "\frac{1}{(D-3D')^2} (12x^2 + 36 xy)"

"= \frac{1}{D^2} \left(1-\frac{3D'}D{}\right)^{-2} (12x^2+36xy)"

"= \left(1+2\frac{3D'}{D}+3\right(\frac{3D'}{D}\left)^2+4\right(\frac{3D'}{D}\left)^3+…\right) \int\int(12x^2+36xy) dxdx"

"= \left(1+6\frac{D'}{D}+27\right(\frac{D'}{D}\left)^2+108\right(\frac{D'}{D}\left)^3+…\right) \left[ x^4 + 6 x^3 y 
 \right]"

"= x^4 + 6 x^3 y + \frac{6}{D} (6x^3)"

"= x^4 + 6x^3 y + 9 x^4 \\ = 10 x^4 + 6x^3 y"

Hence, solution of given differential equation is

"z" = C.F. + P.I.

"\implies z = \phi_1(y+3x)+x\phi_2(y+3x) + 10 x^4 + 6x^3 y"